OpenStudy (anonymous):

When the temperature is 68°F, the speed of sound in air is approximately 343 meters per second. Approximately how many kilometers does sound travel in five seconds?

OpenStudy (mathstudent55):

Multiply 343 m by 5 and convert to km.

OpenStudy (janu16):

The equation is VELOCITY = DISTANCE/TIME. Or V = d/t To solve this equation for d, multiply both sides by t. v*t = d. Now plug in the values for v and t. 343 m/s * 5 s = d 1715 m = d. This answer is in meters. Since there are 1000 m in a kilometer, divide 1715 by 1000 to get km. 1715m * 1 km/1000m = 1.715 km

OpenStudy (janu16):

@KillerMustache2